# Excision Score: Evaluating Edits with Surgical Precision

### v1

@@ Abstract
Many tasks revolve around editing a document, whether code or text. We formulate the revision similarity problem to unify a wide range of machine learning evaluation problems whose goal is to assess a revision to an existing document. This definition rests on the observation that revisions usually change only a small portion of an existing document, so the existing document and its immediate revisions share a majority of their content.
@@ 1 Introduction
These three properties rule out dynamic measures, notably pass@$k$, that rely on execution. Their executability constraint is crippling. Even ignoring NLP tasks, many AI4Code tasks do not produce executable code, like code summarization and commit message generation. Even executable code can be nontestable Weyuker (1982). Even considering only code generation, their utility falters in the face of incomplete codebases. Even restricted to tasks that produce testable code, dynamic measures under-approximate program behaviour Dijkstra (1972), which undermines their interpretability, and can be prohibitively computationally expensive. For example, Neubig & Wang (2024) report that evaluation on some 300 samples of SWE-Bench-like dataset took them 2 days; Adamczewski (2025) managed to reduce it to 1 hour per 500 dataset samples with powerful hardware and dedicated containerized environments optimized for the task. [^1] Although these estimates include the time needed to run the inference of an LLM, we believe they illustrate well the hardships connected with execution-based measures. Thus, an effective measure should be static.
@@
In this work, we proposing the umbrella term “revision similarity” to unite all ML tasks that can be evaluated by three sequences, an original document and two revisions of it, one a golden reference, and the other, the hypothesis to evaluate. We specify five adequacy criteria that any measure of revision similarity should meet and show how many popular measures fail to meet them. We introduce a new measure — ExcisionScore (ES) — that does. It is a static, task-agnostic, interpretable, and lightweight measure. ES aligns a pair of revisions with their source document to focus on their divergent regions, whose $n$-grams it compares.
@@ 3.2 Adequacy Criteria for Revision Similarity
Finally, Property 5 introduces dependence on semantics of the origin document.
Property 5 (Reward Semantically Equivalent Mismatches).
Edits should be permitted to differ as long as they result in semantically equivalent revisions. Often, penalizing lexical differences that result in no semantic difference is seen as undesirable (according to measurable human preference on these tasks).
As a trivial example, consider the case when $O$ contains 3 blank lines, $A$ inserts some text after the first blank line, while $B$ inserts this same text after the second. Although the insertions do not perfectly match, whitespace-only difference might be permissible. As another example, consider a code editing problem where the task is to add a certain method to a class. In most programming languages, method definition order plays no role. Thus, even if the content of the inserted method was correct, it is likely to be penalized for inserting it at the wrong position. In NLP, this Property˜5 is often partially addressed by providing multiple ground truth references in the hope of covering multiple semantic equivalence classes. Thus, mismatch in case of semantic equivalence can be tolerated by the reference set or the measure itself, with the latter approach having some advantages.
Satisfying Property˜5 is undecidable. There are many ways to partially satisfy it. Our new measure, ExcisionScore does so by using n-grams to be locally insensitive to positional mismatches, as we show in Section˜3.4. Empirically, this accounts for many semantically equivalent mismatches in our data sets.
@@ 4 Evaluation: ExcisionScore as Execution Proxy
Quality measures on coding tasks can be categorized into semantic correctness of the produced code and its syntactic similarity to some ground truth reference. Semantic correctness, code’s functional properties are usually assessed by executing a suite of tests shipped with the evaluation dataset. Equipping a dataset with a high-coverage test suite is typically a difficult and expensive task, and so is executing these tests in the specialized environment. Syntactic similarity, on the other hand, is very easy to estimate with metrics such as BLEU/ROUGE/chrF, and the required ground truth label is often cheap to mine by masking out the already existing code. It’s a questionable but widely accepted practice to use syntactic similarity as a cheap proxy to the expensive semantic correctness verification for AI-generated programs. To address this practice, we explore how well ExcisionScore, along with other popular static measures, correlates with test execution.
@@ 5 Related Work
Numerous measures exist to assess the similarity of two strings in different contexts. We argue that all of them are fundamentally ill-suited for the revision similarity problem, as the third string—the origin OO—must be taken into account. Without it, pairwise measures cannot distinguish between revision similarity due to inheriting parts of $O$ unchanged and that due to performing the same edits to $O$. As a result, pairwise measures reward shared context as matching (Section˜2). Pairwise N-gram-based lexical measures include BLEU Papineni et al. (2002), CodeBLEU Ren et al. (2020), CrystalBLEU Eghbali & Pradel (2023), METEOR Banerjee & Lavie (2005), ROUGE Lin (2004), and chrF Popović (2015). BLEU has well-documented limitations, including its inability to address revision similarity — a gap we rigorously analyze in Section˜2. For a systematic critique of BLEU’s shortcomings (e.g., its insensitivity to paraphrasing and granular edits), we direct readers to Callison-Burch et al. (2006) and Reiter (2018). Despite these flaws, BLEU persists as a de facto standard due to its simplicity, reproducibility, and historical inertia.
Evaluating Grammatical Error Correction (GEC) techniques can be seen as an instance of the revision similarity problem. In GEC, edits are typically small and evaluation requires strict measures that are sensitive to word order. Alignment has been used to leverage these aspects in designing GEC metrics such as I-Measure Felice & Briscoe (2015) and $M^{2}$ (MaxMatch) Dahlmeier & Ng (2012). Evaluation of Text Simplification (TS) can, in some cases, also be viewed as revision similarity problem, provided that the simplifying changes do not rewrite the entire text. SARI Xu et al. (2016) (defined in Equation˜2) is an n-gram-based metric designed for the text simplification problem.

### v2

@@ Abstract
Many tasks revolve around editing a document, whether code or text. We formulate the revision similarity problem to unify a wide range of machine learning evaluation problems whose goal is to assess a revision to an existing document. We observe that revisions usually change only a small portion of an existing document, so the existing document and its immediate revisions share a majority of their content.
@@ 1 Introduction
These three properties rule out dynamic measures, notably pass@$k$, that rely on execution. Their executability constraint is crippling. Even ignoring NLP tasks, many AI4Code tasks do not produce executable code, like code summarization and commit message generation. Even executable code can be nontestable Weyuker (1982). Even considering only code generation, their utility falters in the face of incomplete codebases. Even restricted to tasks that produce testable code, dynamic measures under-approximate program behaviour Dijkstra (1972), which undermines their interpretability, and can be prohibitively computationally expensive. For example, Neubig & Wang (2024) report that evaluation on some 300 samples of SWE-Bench-like dataset took them 2 days; Adamczewski (2025) managed to reduce it to 1 hour per 500 dataset samples with powerful hardware and dedicated containerized environments optimized for the task. [^1] Although these estimates include the time needed to run the inference of an LLM, we believe they exemplify the hardships connected with execution-based measures. Thus, a static measure is an indispensable part of the evaluator’s toolbox, necessary for scenarios where dynamic evaluation is infeasible or incomplete.
@@
In this work, we proposing the umbrella term “revision similarity” to unite all ML tasks that can be evaluated by three sequences, an original document and two revisions of it, one a golden reference, and the other, the hypothesis to evaluate. We specify five adequacy criteria that any measure of revision similarity should meet and show how many popular measures fail to meet them. We introduce a new measure — ExcisionScore (ES) — that does. It is a static, task-agnostic, interpretable, and lightweight measure. ES aligns a candidate and a set of reference revisions with their source document to focus on their divergent regions, whose $n$-grams it compares. By constructing its under-approximation from a set of references, ES captures a different subspace of program behavior than dynamic measures, a semantic variability we formalize and explore in Section˜3.2.
@@ 3.2 Adequacy Criteria for Revision Similarity
Finally, Property˜5 introduces dependence on the semantics of the origin document.
Standard reference-based evaluation relies on a set of references $\mathcal{A}$ to define the space of semantically acceptable revisions for a given task, a well-established practice in fields like Grammatical Error Correction (GEC) to account for valid alternative outputs (Bryant et al., 2017, §6). For instance, to evaluate a revision that adds a sort function, $\mathcal{A}$ would contain different correct sorting algorithms (e.g., quicksort, mergesort), thereby defining a range of semantic solutions deemed acceptable.
However, a single semantic solution can often be implemented with minor syntactic variances that do not alter its meaning. Examples include moving a function definition to a different, syntactically valid location within a code file, or applying semantics-preserving code transformations such as inlining (replacing a function call with its body) or outlining (extracting a code segment into a new function). In text, analogous examples include adding a new step before any of its dependencies or changing the order of a bulleted list.
Requiring $\mathcal{A}$ to explicitly enumerate all possible variants of every acceptable solution is computationally expensive. For instance, to handle all possible locations for a single function definition across $\mathcal{O}(|O|)$ locations, one would need a reference for each. Comparing a candidate revision against this set involves a quadratic $\mathcal{O}(|O|^{2})$ number of pairwise block comparisons to align the candidate’s function with each reference’s function. This cost is incurred for each such syntactic variance.
Therefore, a robust similarity measure should be efficient with respect to the reference set $\mathcal{A}$ for such semantics-preserving, syntactic variances. Property˜5 requires that the measure, leveraging knowledge of the language and task, can recognize a candidate revision $B$ as matching a reference $A\in\mathcal{A}$ even if $B$ and $A$ differ only by a variance of this kind, without requiring $\mathcal{A}$ to explicitly enumerate all possible variants. Moreover, the measure should acknowledge a partial match in case some of the edits $B$ makes achieve the effect semantically equivalent to some of $A$’s edits, even if $A$ otherwise semantically differs from $B$.
Example 4 (Tolerated Syntactic Variances).
In the sort function task, the set $\mathcal{A}$ defines the acceptable semantic range (e.g., quicksort, mergesort). Property˜5 ensures that if a candidate revision $B$ places a correct quicksort implementation in a different location than the reference quicksort in $\mathcal{A}$, it is still correctly identified as matching the quicksort semantics. This avoids the need for a separate reference for every possible function location, making the evaluation both practical and semantically grounded.
To formalize the notion of semantics-preserving syntactic variances, we first define a semantic equivalence relation between documents under a set of semantics-preserving transformations.
Definition 3 (Equivalence under Semantics-Preserving Transformations).
Let $\llbracket s\rrbracket_{L}$ denote the semantics of a string $s$ in language $L$. Let $\equiv_{L}$ be a semantics-preserving syntactic equivalence relation on strings in $L$, such that $A\equiv_{L}B$ implies $\llbracket A\rrbracket_{L}=\llbracket B\rrbracket_{L}$. This relation is defined by a set of syntactic transformation rules (e.g., function relocation, inlining) that are known a priori to preserve semantics. Any practical measure uses a concrete $\equiv_{L}$ that under-approximates the full, undecidable semantic equivalence relation.
We now extend this concept to define semantic equivalence between sets of edits.
An atomic edit is a tuple defining a single, irreducible change to a sequence, e.g., of tokens or characters. It combines an operation (insert, delete, replace, swap, etc.), a target location index within the sequence, and an optional operand: a new element for insert/replace, an index for swap, ignored for delete. Let $A$ be the revision produced by applying a set of atomic, nonoverlapping edits $\{a_{1},\dots,a_{n}\}$ to $O$, denoted $A=\{a_{1},\dots,a_{n}\}(O)$. Similarly, let $B=\{b_{1},\dots,b_{m}\}(O)$. By nonoverlapping, we mean that the indices of distinct edits are distinct: formally, for $a_{i}=(\_,x,\_),a_{j}=(\_,y,\_)\in\{a_{1},\dots,a_{n}\},i\neq j\Rightarrow x\neq y$, where ’_\_’ denotes don’t care. The edits in a nonoverlapping set can be applied simultaneously. For any subset of indices $I\subseteq\{1,\dots,n\}$, we denote by $a_{I}(O)=\{a_{i}\}_{i\in I}(O)$ the revision obtained by applying exactly those edits to $O$. For example, $\{\}(O)=O$ and $\{a_{i}\}_{i=1}^{n}(O)=A$.
Definition 4 (Semantically Equivalent Edits).
For $\{a_{i}\}_{1}^{n}$ and $\{b_{i}\}_{1}^{m}$, let $I\subseteq 1..n$, $J\subseteq 1..m$, $\bar{I}=1..n\setminus I$, $\bar{J}=1..m\setminus J$. Assume that the indices of the edits $b_{J}$ do not overlap with the indices of $a_{\bar{I}}$ and indices of $a_{I}$ do not overlap with those of $b_{\bar{J}}$.
The subsets of edits $a_{I}=\{a_{i}\mid i\in I\}$ and $b_{J}=\{b_{j}\mid j\in J\}$ are semantically equivalent under $\equiv_{L}$ iff:
$A\equiv_{L}(b_{J}\cup a_{\bar{I}})(O)\quad\land\quad B\equiv_{L}(a_{I}\cup b_{\bar{J}})(O)$
This relation is symmetric. It formalizes the condition that the edits in $a_{I}$ and $b_{J}$ are interchangeable syntactic variances for achieving the same semantic outcome within the context of their respective revisions; in other words, the edits are equivalent in the surrounding context of the other edits $A$ and $B$ apply. Finding the subsets $I$ and $J$ is, of course, undecidable in general, but, in practice, can be under-approximated with knowledge of the transformations that underlie a particular concretisation of $\equiv_{L}$.
Property 5 (Obliviousness to Semantically Equivalent Syntactic Variances).
In the notation of Definition 4, a similarity measure $m$ must satisfy:
$m(A,B;O)=m(A,(a_{I}\cup b_{\bar{J}})(O);O),$
whenever the edits $\{a_{i}\mid i\in I\}$ and $\{b_{j}\mid j\in J\}$ are semantically equivalent under $\equiv_{L}$.
This property requires a revision similarity measure to be oblivious to the choice between syntactically different but semantically equivalent edit sets under some realisation of $\equiv_{L}$. Our new measure, ExcisionScore does so by being insensitive to the order of n-grams, as we show in Section˜3.4. This accounts for semantically equivalent mismatches that can be fixed by reordering blocks of content, which we found to be common in our our evaluation datasets.
Humans care about more than mere semantics when comparing revisions. For instance, a human may prefer unobfuscated or well-refactored code or, in text, a paraphrase or a summary of some topic. Humans, to take another example, also vary greatly in terms of their code commenting preferences. Property˜5 permits handling these aspects in two ways: either by defining $\equiv_{L}$ to consider them or by including examples of these aspects in the set of references. Using $\equiv_{L}$ to do so effectively makes the relevant aspects semantic.
@@ 4 Evaluation: ExcisionScore as Execution Proxy
Actively developed real-world codebases often include incomplete, non-compilable code for which tests have not yet been written. Even for compilable code in large systems, full build and test cycles can be prohibitively long, making rapid, lightweight static feedback essential. Equipping a dataset with extensive test coverage can be more difficult than mining ground truth solutions. For these reasons, static measures based on syntactic similarity to the reference solution persist as cheap proxies to expensive verification for AI-generated programs. Following widely accepted practice, we therefore explore how well ExcisionScore and other popular static measures correlate with test execution. Our evaluation approach answers the question: "When execution is possible, which static metric best predicts its outcome?".
@@ Results
We found that 2% of LLM solutions for HumanEvalFix dataset and 5% for CanItEdit match the reference solution exactly. Excluding them from the data decreases all the correlation coefficients by 3-9%, leaving the ranking of different metrics unaffected on HumanEvalFix. Differences in correlation coefficients remain insignificant on CanItEdit.
@@ 5 Related Work
Numerous measures assess the similarity of two strings in different contexts. We argue that they are ill-suited for the revision similarity problem, because they do not take the original document $O$ string into account. Without $O$, pairwise measures cannot distinguish between revision similarity due to inheriting parts of $O$ unchanged and that due to performing the same edits to $O$. As a result, pairwise measures reward shared context. Pairwise N-gram-based lexical measures include BLEU Papineni et al. (2002), METEOR Banerjee & Lavie (2005), ROUGE Lin (2004), and chrF Popović (2015). BLEU has well-documented limitations, including its inability to address revision similarity — a gap we rigorously analyze in Section˜2. For a systematic critique of BLEU’s shortcomings (e.g., its insensitivity to paraphrasing and granular edits), we direct readers to Callison-Burch et al. (2006) and Reiter (2018). Despite these flaws, BLEU persists as a de facto standard due to its simplicity, reproducibility, and historical inertia. Syntax-aware pairwise static measures for code include Tree Edit Distance Schwarz et al. (2017) (TED), RUBY Tran et al. (2019), CodeBLEU Ren et al. (2020), and CrystalBLEU Eghbali & Pradel (2023). They require the code to be parseable. TED is very interpretable but takes $\mathcal{O}(n^{3})$ time, where $n$ is the number of nodes in the syntax tree. CrystalBLEU removes ‘shared’ n-grams, but interprets ‘shared’ as frequent n-grams in the language, not unedited tokens inherited from $O$ as in our work. Lastly, embedding-based measures such as cosine similarity, BERTScore Zhang et al. (2020) and CodeBERTScore Zhou et al. (2023), while capturing semantic similarity, remain fundamentally pairwise and cannot account for the original document $O$.
Evaluating Grammatical Error Correction (GEC) techniques can be seen as an instance of the revision similarity problem. In GEC, edits are typically small and evaluation requires strict measures that are sensitive to word order. Alignment has been used to leverage these aspects in designing GEC metrics such as I-Measure Felice & Briscoe (2015) and $M^{2}$ (MaxMatch) Dahlmeier & Ng (2012). Evaluation of Text Simplification (TS) can, in some cases, also be viewed as revision similarity problem, provided that the simplifying changes do not rewrite the entire text. SARI Xu et al. (2016) (defined in Equation˜2) is an n-gram-based metric designed for text simplification.